# Long Context and Long Form Output Dataset and Benchmarking

### v2

@@ 2.7 Data Statistics
BookSum 27 Litertyre: books, novel, act 4114
@@
GovReport 25 Political: Congressional Research Service and
US Government Accountability Office 5078
@@
Table 2: LCFO Summary: Domains and Statistics
@@ 3 Experiments
We experimented with closed and open LLMs. We chose GPT-4o-mini-64k [^1] https://openai.com/index/gpt-4o-mini-advancing-cost-efficient-intelligence/ for the closed model and Llama-3.1-70B (Dubey et al., 2024) for the open-source one. For summarization, we ran the model with all length ratios (5%, 10%, 20%), while for summary expansion, we only expanded the summaries 20% to the full document. We also performed a postprocessing step to filter the templated response such as ‘‘**Summary**’’, ‘‘Here is the summary:’’, etc.
@@
Table 3 shows the general results of the selected models at different levels of gradual summarization. Results broken down by domains are reported in the Appendix H. Note that LLMs tend to perform similarily regardless of the length of the output in terms of human scores. This is not the case for humans that show to lag behind when performing short summaries.
@@
The best results are consistently achieved with GPT-4o-mini and are consistent with previous research findings (Que et al., 2024). This model even surpasses human-level quality in short summaries. This may be explained by humans tending to perform worse when summarizing short documents and better when summarizing long ones.
@@
Table 4 shows the overall results on the summary expansion task only on a factor of 5. The performance of models is not coherent across metrics that look only at the output (i.e.%WC, REP-3, CoLA, COH-2, and AVG). In terms of HE and coherently with the human evaluation results, GPT-4o-mini is the best performing model.
We do not report results on expanding by larger factors (10 or 20) since models are performing poorly (see some examples in the Appendix H). The main limitation in this case is that the models do not reach the required output length.
@@
When comparing across tasks (i.e. summarizing to 20% or doing summary expansion from 20% by a factor of 5), the results show better performance in the former. It is expected that summary expansion is a more challenging task across domains and all models. Current models struggle with this task. If we compare HE, the deltas in the same model vary from 6% for GPT-4o-mini to $\approx$ 30% for Llama-3.1-70B. When comparing output-based metrics, there are discrepancies in conclusions (i.e., Llama-3.1-8B better than 70B model). However, HE is still worse for the 8B model. This may indicate that selected output-based quality metrics are less reliable than the HE score (see the analysis below for metrics evaluation).
@@
Table 5: Spearman correlation coefficients (and p-value) for various aspects and overall scores between automatic metrics and human evaluation for summarization (S) and summary expansion (SE). For the former, we show correlations between CoLA and Human evaluation Q2d; SH-4 and Human evaluation Q2a; SH-5 and Human evaluation Q2b and R-L/AVG/HE and Human evaluation Q3 (appendix E). For the latter, we show correlations between AVG/HE and Human evaluation Q3 (appendix F).
@@
In our study, we consider human evaluation, conducted according to the guidelines outlined in Appendices E and F, as the definitive measure of the overall quality score, as well as the scores for individual quality aspects such as coverage and attribution. To mitigate potential biases among the annotators, we calculate the average of three annotations for each task.
Table 5 presents the Spearman correlation coefficients for various aspects and overall scores, comparing automatic metrics and human evaluations for summarization and summary expansion, respectively. A higher Spearman correlation coefficient signifies a stronger correlation between the automatic metrics and human annotation. The metric that shows the highest correlation with human annotation corresponds to SH-4, which measures attribution. When comparing metrics that measure overall performance, we observe that R-L is not very good at correlation, but it may also be due to the fact that our task is not the best suited to use human references. HE is the one with the highest correlation. AVG low-correlation (both in S and SE) may be explained by the fact that individual averaged metrics are not very good or they cover more specific aspects which may not end capturing the overall performance. This low correlation for R-L and AVG can explain the discrepancy observed in the model ranking (specially between Llama-3.1-70B and 8B in Tables 3 and 4. Correlations are low in all cases, which shows the difficulty of the evaluation. Beyond the challenge of automatizing it, we should add the fact that humans struggle in generating short summaries, which may imply that humans also struggle in evaluating them.
@@ 4 Related Work
Related work on long context and long form output comes in many flavors. In this section, we cover a summary of related works on long context and long-form output benchmarking as well as a special focus on summarization evaluation.
@@
Infinite length datasets such as NIAH, RULER (Hsieh et al., 2024) work with distracting information. Finite-length nondistractive-based datasets include: Longbench (Bai et al., 2024) and Marathon (Zhang et al., 2024b) that includes tasks with 5–25k context and, more recently, Kwan et al. (2024) build a dataset up to 8k tokens context length to evaluate LLMs’ long-context understanding across five key abilities: understanding of single or multiple relevant spans in long contexts based on explicit or semantic hints, and global context understanding. Loong is a multi-document QA dataset up to 200k context to assess RAG abilities. HelloBench (Que et al., 2024) includes summarization of a selection of long-input documents (3k to 6k word length).
@@
There is a lack of reference-based datasets on long form output. However, there are datasets that study prompting of different long-form generation; e.g., StoryGen (Zhu et al., 2023) includes prompts to generate stories and HelloBench is one of the most diverse text long form generation including stories, screenplays, keyword writing.
Summarization metrics.
Most popular metrics are string-based, such as ROUGE or BLEU or meaning-representation score. More complex meaning-representation scores (e.g., BERTScore, Zhang et al., 2020) and/or supervised-metrics have been trained with success on human annotations of several benchmarks; e.g., Seahorse (Clark et al., 2023), summeval (Fabbri et al., 2021), realsumm (Bhandari et al., 2020). Several alternative to using summaries as gold references include the Atomic Content Units (ACU) (Liu et al., 2023a), which is based on fine-grained semantic units and allows for a high inter-annotator agreement; question-based summaries (QMSum, sQuality) or human QA-based methods (based on human pre-annotated QA) which were introduced by Clarke and Lapata (2010) for complementing string-based metrics, culminating in one of the largest datasets with NarrativeQA (46k questions). Recently, QA-based approaches have followed an automatic approach instead of human annotation to generate QA (QuestEval, Scialom et al., 2021). On human evaluation, there are several relevant works among which we find the Seahorse guidelines (Clark et al., 2023). Some hints (Krishna et al., 2023) on the task are (1) FINE-grained annotations have lower inter-annotator variance than COARSE-grained annotations; (2) partially annotating a summary reduces annotator workload while maintaining accuracy; (3) highlighting hints in the source document has limited usefulness for evaluating long-form summaries.
Long-form metrics.
Variety of criteria can be found in Sun et al. (2022) and Gao et al. (2024) based on a combination of fluency, diversity and semantics. Other types of evaluation focus more on coherence (Goyal et al., 2022) and factuality (Song et al., 2024)). LLMs have been used as evaluators in works like Chiang and Lee (2023); Liu et al. (2023b). While optimistic, findings show room for improvement in verification and consistency (Shen et al., 2023). QA-based methods with human annotations are used in Tan et al. (2024) focusing on informativeness and coverage of the text output. On the human evaluation, a related work are the StoryGen guidelines (Zhu et al., 2023), among others, showing the difficulty of the task,
Our contribution on datasets involve the manual collection of 3-length summaries from long input documents. This collection also includes abstractive QA (non-multiple choice) to test comprehension. Our contribution on metrics involves new human evaluation protocols on summarization and summary expansion, as well as annotations to train/evaluate supervised metrics on long-form outputs.
@@ Appendix B Data Details
Appendix C Summarization Guidelines
@@
Appendix D Human Summary Sample
@@
D.1 Source Document Excerpt
@@
D.2 Long, Medium and Short Summaries
@@ D.3 Question and Answer Set
Appendix E Summarization Human Evaluation Guidelines
@@
Appendix F Summary Expansion Human Evaluation Guidelines
@@
Appendix G Prompting Details
The prompts contain three parts: General guideline, domain-specific prompts, and input context. The general guideline adapts the human guidelines (Appendix C) for the summarization and summary expansion, while the domain specific prompts give extra information about the domain as instructions of expected output. In the prompt template below, the general guideline is provided, {{domain-X}} denotes the domain-specific prompt. {{input}} is for the input document for the summarization task and human summaries for the summary expansion task.
@@ Appendix G Prompting Details
The model-specific prompts for each domain is listed below. Note that not all domains have the prompt template for summary expansion.
@@
    Appendix H Detailed Results
    Table 6 shows the detailed results of different models in the summary expansion task. It is shown that, despite having repeated instructions on the length, all models behave greatly differently in different domains. In particular, GPT-4o-mini can generate longer scientific/technical texts, but struggle to generate longer texts in conversational texts without sacrifying the qualities. On the other hand, medium-sized models such as Llama-3.1-8B generate more texts consistently across domains, but at a higher repetition.
@@ Appendix H Detailed Results
    Llama-3.1-70B 1.449 0.782 0.814 0.500 0.386 41.466 3.800
    Llama-3.1-8B 1.495 0.828 0.851 0.568 0.418 36.493 4.480
    Table 6: Performance on the summary expansion task per dataset.
    The results of the summarization task of different levels are detailed in tables 78 and 9. According to human evaluation, the best performing result in Table 7 is with GPT-4o-mini in the wikipedia domain and in Table 8 is with the same model in the legal domain (LexGlue). For table 9, the best results are with human output in the conversational domain (Summscreen).
@@
    Llama-3.1-70B 0.417 0.491 0.963 0.831 0.660 0.480 0.489 60.241 6.440
    Llama-3.1-8B 0.402 0.474 0.968 0.839 0.642 0.476 0.490 60.415 7.080
    Table 8: Performance on the 10-percent summarization task per dataset
@@
    Llama-3.1-70B 0.454 0.589 0.957 0.834 0.615 0.441 0.452 54.075 6.600
    Llama-3.1-8B 0.469 0.642 0.960 0.847 0.601 0.456 0.444 54.142 6.760
    Table 9: Performance on the 20-percent summarization task per dataset.

### v3

@@ 2.7 Data Statistics
BookSum 27 Literature: books, novel, action 4114
@@
GovReport 25 Political: Congressional Research Service and US Government Accountability Office 5078
@@
Table 1: LCFO Summary: domains and statistics (number of documents N and average length in words len)
@@ 3 Experiments
We experimented with closed and open LLMs. We chose GPT-4o-mini-64k [^4] https://openai.com/index/gpt-4o-mini-advancing-cost-efficient-intelligence/ for the closed model and LLaMA 3.1-70B (Dubey et al., 2024) for the open-source one. For summarization, we ran the model with all length ratios (5%, 10%, 20%), while for summary expansion, we only expanded the summaries 20% to the full document. We also performed a postprocessing step to filter the templated response such as ‘‘**Summary**’’, ‘‘Here is the summary:’’, etc.
@@
Table 2 shows the general results of the selected models at different levels of gradual summarization. Results broken down by domains are reported in the Appendix J. Note that LLMs tend to perform similarly regardless of the length of the output in terms of human scores. This is not the case for humans that show to lag behind when performing short summaries. The best results are consistently achieved with GPT-4o-mini and are consistent with previous research findings (Que et al., 2024). This model even surpasses human-level quality in short summaries. This may be explained by humans tending to perform worse when summarizing short documents and better when summarizing long ones.
@@
Table 3 shows the overall results on the summary expansion task by a factor of 5, giving the 20% summary input. The performance of models is not coherent across metrics that look only at the output (i.e.%WC, REP-3, CoLA, COH-2, and AVG). In terms of HE and coherently with the human evaluation results, GPT-4o-mini is the best performing model. Additionally, we report the results of other combinations, for example, expanding by larger factors (10 and 20) giving 10% summary input and giving 5% summary input, respectively, in Appendix J.
@@
When comparing across tasks (i.e. summarizing to 20% or doing summary expansion from 20% by a factor of 5), the results show better performance in the former. It is expected that summary expansion is a more challenging task across domains and all models. Current models struggle with this task. If we compare HE, the deltas in the same model vary from 6% for GPT-4o-mini to $\approx$ 30% for LLaMA 3.1-70B. When comparing output-based metrics, there are discrepancies in conclusions (i.e., LLaMA 3.1-8B better than 70B model). However, HE is still worse for the 8B model. This may indicate that selected output-based quality metrics are less reliable than the HE score (see the analysis below for metrics evaluation).
@@
Table 4: Spearman correlation coefficients (and p-value) for various aspects and overall scores between automatic metrics and human evaluation for summarization (S) and summary expansion (SE). For the former, we show correlations between CoLA and Human evaluation Q2d; SH-4 and Human evaluation Q2a; SH-5 and Human evaluation Q2b and R-L/AVG/HE and Human evaluation Q3 (appendix G). For the latter, we show correlations between AVG/HE and Human evaluation Q3 (appendix H).
@@
In our study, we consider human evaluation, conducted according to the guidelines outlined in Appendices G and H, as the definitive measure of the overall quality score, as well as the scores for individual quality aspects such as coverage and attribution. To mitigate potential biases among the annotators, we calculate the average of three annotations for each task. Table 4 presents the Spearman correlation coefficients for various aspects and overall scores, comparing automatic metrics and human evaluations for summarization and summary expansion, respectively. A higher Spearman correlation coefficient signifies a stronger correlation between the automatic metrics and human annotation. The metric that shows the highest correlation with human annotation corresponds to SH-4, which measures attribution. When comparing metrics that measure overall performance, we observe that R-L is not very good at correlation, but it may also be due to the fact that our task is not the best suited to use human references. HE is the one with the highest correlation. AVG low-correlation (both in S and SE) may be explained by the fact that individual averaged metrics are not very good or they cover more specific aspects which may not end capturing the overall performance. This low correlation for R-L and AVG can explain the discrepancy observed in the model ranking (specially between LLaMA 3.1-70B and 8B in Tables 2 and 3. Correlations are low in all cases, which shows the difficulty of the evaluation. Beyond the challenge of automatizing it, we should add the fact that humans struggle in generating short summaries, which may imply that humans also struggle in evaluating them.
@@ 4 Related Work
Related work on long context and long form output comes in many flavors. We cover a summary on long context and long-form output datasets.
@@
Infinite length datasets such as NIAH, RULER (Hsieh et al., 2024) work with distracting information. Finite-length nondistractive-based datasets include: Multi-LexSum (Shen et al., 2022) is a collection of 9,280 expert-authored summaries drawn from single domain (Civil Rights Litigation Clearinghouse) writing the length of the source documents often exceeds two hundred pages per case and summaries are presented in two-sizes: a short and longer version; Longbench (Bai et al., 2024) and Marathon (Zhang et al., 2024b) that includes tasks with 5–25k context and, more recently, Kwan et al. (2024) build a dataset up to 8k tokens context length to evaluate LLMs’ long-context understanding across five key abilities: understanding of single or multiple relevant spans in long contexts based on explicit or semantic hints, and global context understanding. Loong is a multi-document QA dataset up to 200k context to assess RAG abilities. HelloBench (Que et al., 2024) includes summarization of a selection of long-input documents (3k to 6k word length).
@@
There is a lack of reference-based datasets on long form output. However, there are datasets that study prompting of different long-form generation; e.g., StoryGen (Zhu et al., 2023) includes prompts to generate stories, and HelloBench is one of the most diverse long form generation benchmarks including stories, screenplays, keyword writing.
Our contribution on datasets involve the manual collection of 3-length summaries from long input documents. This collection also includes abstractive QA (non-multiple choice) to test comprehension. Our contribution on metrics involves new human evaluation protocols on summarization and summary expansion, as well as annotations to develop supervised metrics on long-form outputs.
@@ 5 Conclusions
Computing
In terms of computing, evaluating LLMs on LCFO benchmark require larger memory due to both its big context size and the long-form output (should the models be capable to it). In case of LLaMA, we used 1 NVIDIA GPU A100 80 GB for the 8B model, and 8 GPUs for the 70B model. The resource was shared with the loading of scoring models (SH, CoLA) as well. Each evaluation run over 10 domains takes 90 minutes, including the computation of all the scores except HelloEval (where the computing time depends on the external availability of GPT-4 endpoint deployment).
Annotations
Annotators were paid a fair rate. Each of the annotators signed a consent form agreeing on the dataset and its usage that they were participating in.
@@ Appendix B Data Details
Appendix C Summary evaluation
Table 5 summarises the metrics used to evaluate.
Task Area Metric Description Reference
Sum Attribution SH-4 Seahorse-Large-Q4 score Clark et al. (2023)
SumExp Word Count WC
Sum/SumExp Overall AVG Empirical average of metrics
Sum/SumExp Overall HE HelloEval score Que et al. (2024)
Table 5: Summary of automatic metrics used in different tasks.
Appendix D Selection of automatic metrics
At an early stage of our work (before collecting human annotations), we came up with a list of candidate metrics of summarization quality. We provide it in Table 6.
To select a single metric per aspect, we analyzed their Spearman correlations with human annotations of the English subset of SEAHORSE (Clark et al., 2023), focusing on Questions 2 to 6: absence of repetitions, grammatical correctness, attribution to the source, coverage of the source, and conciseness (which also reflects overall quality). We did not work with Question 1 (comprehensibility), because our metrics focus on inter-sentential comprehensibility, whereas 66% of SEAHORSE summaries consist of a single sentence.
We report the resulting correlations, grouped by the three English SEAHORSE subsets, in Table 7. One can see that the metrics we selected for summarization correlate reasonably with their corresponding aspects across all 3 subsets most of the time [^7] An exception is the grammaticality/fluency aspect (Question 2) on the Wikihow subset, where no metric demonstrated adequate correlation with human annotations., which justifies our choice of them.
Metric Short name Implementation
ref_rouge_fmeasure R-L ROUGE-L F1 score, computed with the rouge-score Python package (in the rougeLsum mode)
ref_rouge_recall ROUGE-L recall (see above)
ref_rouge_precision ROUGE-L precision (see above)
char_len_ratio Summary-to-source lengths ratio, in characters
sent_len_ratio Summary-to-source lengths ratio, in sentences
word_len_ratio Summary-to-source lengths ratio, in words
td_coherence_1st 1st order coherence, computed with the textdescriptives Python package
td_coherence_2nd COH-2 2nd order coherence, computed with the textdescriptives Python package
td_flesch_kincaid_grade Flesch-Kincaid grade readability score (textdescriptives)
td_gunning_fog Gunning-Fog readability score (textdescriptives)
td_dependency_distance_mean Mean distance of syntactic dependencies (textdescriptives)
word_ngram_src_overlap Fraction of source word n-grams (1 to 3) that appear in the summary
word_ngram_repetition_rate REP-3 Fraction of summary word n-grams (1 to 3) that appear more than once
src_rouge_recall ROUGE-2 recall of the summary w.r.t. the source
src_rouge_f1 ROUGE-2 F1 score of the summary w.r.t. the source
src_rouge_precision ROUGE-2 precision of the summary w.r.t. the source
tgt_fluency CoLA Mean predicted probability of the sentence being linguistically acceptable, as per the model from Krishna et al. (2020)
src_fluency_diff Difference of mean predicted probabilities of being acceptable in the summary and in the source
mean_self_sonar_sim Mean cosine similarity SONAR embeddings of each summary sentence to its most similar summary sentence
mean_src_sonar_sim Mean cosine similarity SONAR embeddings of each summary sentence to its most similar source sentence
mean_src_coverage_sonar_sim Mean cosine similarity SONAR embeddings of each source sentence to its most similar summary sentence
mean_monotonic_src_sonar_sim Mean cosine similarity SONAR embeddings of each summary sentence to its monotonically aligned source sentence
sonar_sim_monotonicity Ratio of mean_monotonic_src_sonar_sim to mean_src_sonar_sim
p_entail_full Probability of the summary being entailed by the source, predicted with tasksource/deberta-small-long-nli model
p_noncontradict_full Probability of the summary not contradicting the source, predicted with tasksource/deberta-small-long-nli model
smollm_reconstruct_loss Cross-entropy loss of the source given the summary, computed with the HuggingFaceTB/SmolLM-360M-Instruct model
seahorse_q4 SH-4 Probability of the “Yes” response extracted from the SEAHORSE Q4 model (google/seahorse-large-q4)
seahorse_q5 SH-5 Probability of the “Yes” response extracted from the SEAHORSE Q5 model
seahorse_q6 Probability of the “Yes” response extracted from the SEAHORSE Q6 model
Table 6: The automatic metrics we considered as candidates, along with their implementation details.
SEAHORSE subset wiki xlsum xsum
Quality aspect q2 q3 q4 q5 q6 q2 q3 q4 q5 q6 q2 q3 q4 q5 q6
ref_rouge_fmeasure (R-L) 0.33 0.03 0.11 0.25 0.27 0.17 0.12 0.12 0.23 0.18 0.11 0.11 0.13 0.16 0.15
ref_rouge_recall 0.31 0.02 0.15 0.26 0.25 0.16 0.11 0.10 0.23 0.18 0.11 0.09 0.12 0.17 0.14
ref_rouge_precision 0.29 0.02 0.03 0.16 0.20 0.17 0.11 0.14 0.20 0.18 0.12 0.10 0.14 0.14 0.15
char_len_ratio 0.03 0.08 -0.04 -0.04 -0.06 0.00 0.03 0.09 0.20 0.13 0.05 -0.01 -0.06 0.12 0.03
sent_len_ratio -0.02 0.14 -0.06 -0.05 -0.07 -0.04 -0.02 0.08 0.13 0.08 0.08 -0.02 -0.05 0.05 0.01
word_len_ratio 0.03 0.08 -0.05 -0.05 -0.07 -0.02 0.03 0.07 0.18 0.11 0.06 -0.02 -0.07 0.11 0.02
td_flesch_kincaid_grade 0.08 -0.01 0.10 0.11 0.11 0.04 0.07 0.01 0.11 0.09 0.01 0.00 0.10 0.12 0.12
td_gunning_fog 0.14 -0.03 0.12 0.15 0.15 0.02 0.08 -0.00 0.07 0.05 0.05 -0.03 0.10 0.13 0.13
td_dependency_distance_mean 0.07 -0.05 -0.01 -0.01 -0.04 0.09 0.03 -0.02 0.09 0.04 0.06 -0.01 -0.03 0.07 0.05
word_ngram_src_overlap 0.18 0.09 0.08 0.06 0.05 0.06 0.09 0.14 0.18 0.15 0.03 -0.01 0.05 0.15 0.08
word_ngram_repetition_rate (REP-3) -0.59 0.05 -0.13 -0.19 -0.23 -0.31 -0.08 -0.13 -0.05 -0.11 -0.27 -0.05 -0.02 0.04 -0.01
src_rouge_recall 0.06 0.09 0.08 0.06 0.03 0.04 0.05 0.11 0.09 0.11 -0.05 -0.01 0.07 0.13 0.08
src_rouge_f1 0.07 0.10 0.09 0.07 0.04 0.04 0.05 0.11 0.09 0.11 -0.05 -0.01 0.07 0.13 0.08
src_rouge_precision 0.10 0.07 0.20 0.12 0.09 0.05 0.03 0.09 -0.03 0.05 -0.09 -0.00 0.09 0.07 0.06
tgt_fluency (CoLA) 0.07 0.09 0.06 0.08 0.09 0.35 0.25 0.10 0.09 0.16 0.23 0.20 0.18 0.12 0.14
src_fluency_diff 0.06 0.08 0.00 -0.01 0.01 0.29 0.18 0.05 0.06 0.14 0.19 0.21 0.10 0.06 0.08
mean_self_sonar_sim -0.57 0.09 -0.13 -0.14 -0.16 0.02 0.03 -0.02 -0.03 -0.02 0.05 0.02 0.00 0.06 -0.01
mean_src_sonar_sim -0.10 0.06 0.16 0.03 0.02 0.01 0.07 0.08 0.00 0.01 -0.05 0.01 0.09 0.04 0.07
mean_src_coverage_sonar_sim 0.09 0.08 0.10 0.10 0.07 -0.01 0.09 0.11 0.09 0.06 -0.04 0.01 0.10 0.06 0.09
mean_monotonic_src_sonar_sim 0.10 0.03 0.21 0.13 0.11 0.01 0.07 0.08 0.00 0.01 -0.05 0.01 0.09 0.04 0.07
sonar_sim_monotonicity 0.37 -0.07 0.16 0.13 0.10 -0.04 -0.00 0.08 0.05 0.06 -0.02 0.05 -0.04 -0.07 0.02
p_entail_full -0.02 0.02 0.38 0.17 0.20 -0.04 0.00 0.21 0.08 0.14 -0.02 -0.00 0.39 0.26 0.31
p_noncontradict_full -0.02 -0.03 0.20 0.04 0.06 -0.08 -0.05 0.07 -0.03 -0.00 0.01 -0.04 0.25 0.18 0.20
smollm_reconstruct_loss -0.09 -0.05 -0.11 -0.03 -0.03 -0.01 -0.01 -0.06 -0.11 -0.07 0.05 0.02 -0.17 -0.11 -0.13
seahorse_q4 (SH-4) 0.24 0.03 0.52 0.37 0.41 0.19 0.08 0.31 0.17 0.27 0.14 0.06 0.54 0.35 0.43
seahorse_q5 (SH-5) 0.43 0.03 0.33 0.51 0.47 0.20 0.12 0.25 0.27 0.29 0.16 0.14 0.36 0.30 0.31
seahorse_q6 0.45 0.01 0.42 0.51 0.52 0.26 0.10 0.31 0.22 0.31 0.20 0.10 0.47 0.35 0.41
Table 7: Spearman correlations of candidate automatic metrics with human judgments on the English subset of SEAHORSE (Clark et al., 2023), grouped by subsets of the data.
Appendix E Summarization Guidelines
@@
Appendix F Human Summary Sample
@@
F.1 Source Document Excerpt
@@
F.2 Long, Medium and Short Summaries
@@ F.3 Question and Answer Set
Appendix G Summarization Human Evaluation Guidelines
@@
Appendix H Summary Expansion Human Evaluation Guidelines
@@
Appendix I Prompting Details
The prompts contain three parts: General guideline, domain-specific prompts, and input context. The general guideline adapts the human guidelines (Appendix E) for the summarization and summary expansion, while the domain specific prompts give extra information about the domain as instructions of expected output. In the prompt template below, the general guideline is provided, {{domain-X}} denotes the domain-specific prompt. {{input}} is for the input document for the summarization task and human summaries for the summary expansion task.
@@ Appendix I Prompting Details
The model-specific prompts for each domain are listed below. Note that not all domains have the prompt template for summary expansion.
@@
    Appendix J Detailed Results
    Complementing Table 3 from Section 3, Tables 8 and 9 show results of different models in the summary expansion tasks on a different level of summary and expansion. Tables 10, 11 and 12 show the detailed results breaked down by domain on the summary expansion task given the summaries (5%, 10% and 20%) and expanding by a respective factor of (20, 10, 5). It is shown that, despite having repeated instructions on the length, all models behave greatly differently in different domains. In particular, GPT-4o-mini can generate longer scientific/technical texts, but struggle to generate longer texts in conversational texts without sacrifying the qualities. On the other hand, medium-sized models such as LLaMA 3.1-8B generate more texts consistently across domains, but at a higher repetition.
[…]
    GPT-4o-mini 4.789 0.623 0.910 0.564 0.450 67.617
    LLaMA 3.1-70B 2.807 1.830 0.913 0.625 0.391 54.344
    LLaMA 3.1-8B 1.788 0.781 0.906 0.752 0.501 42.572
    Table 8: Performance on the summary expansion task by a factor of 10, given the 10% summary input.
[…]
    GPT-4o-mini 7.960 0.592 0.909 0.574 0.455 65.951
    LLaMA 3.1-70B 2.296 0.631 0.871 0.720 0.488 51.539
    LLaMA 3.1-8B 3.805 0.720 0.967 0.862 0.562 45.141
    Table 9: Performance on the summary expansion task by a factor of 20, given the 5% summary input.
[…]
    BookSum GPT-4o-mini 7.116 0.420 0.953 0.701 0.523 82.421
    LLaMA 3.1-70B 3.307 0.631 0.931 0.863 0.556 66.252
    LLaMA 3.1-8B 5.070 2.170 0.975 0.787 0.442 60.393
    SQuALITY GPT-4o-mini 7.335 0.398 0.954 0.695 0.523 55.620
    LLaMA 3.1-70B 3.010 0.530 0.946 0.769 0.536 47.703
    LLaMA 3.1-8B 3.166 0.624 0.947 0.795 0.539 42.532
    FacetSum GPT-4o-mini 9.348 0.635 0.936 0.593 0.467 59.238
    LLaMA 3.1-70B 0.253 0.792 0.769 0.816 0.475 47.145
    LLaMA 3.1-8B 0.369 0.998 0.843 0.841 0.495 27.660
    Summscreen GPT-4o-mini 8.042 0.914 0.794 0.309 0.307 74.779
    LLaMA 3.1-70B 2.615 0.572 0.836 0.432 0.385 50.914
    LLaMA 3.1-8B 3.370 0.712 0.840 0.474 0.391 44.121
    Table 10: Performance on the summary expansion task by a factor of 20, given the 5% summary input, per dataset.
[…]
    BookSum GPT-4o-mini 5.155 0.443 0.955 0.689 0.518 82.421
    LLaMA 3.1-70B 3.596 1.453 0.981 0.817 0.503 74.280
    LLaMA 3.1-8B 2.451 0.778 0.950 0.859 0.551 60.163
    SQuALITY GPT-4o-mini 3.531 0.418 0.952 0.665 0.511 56.867
    LLaMA 3.1-70B 3.104 1.098 0.957 0.718 0.485 50.855
    LLaMA 3.1-8B 2.056 0.659 0.951 0.813 0.544 38.008
    FacetSum GPT-4o-mini 5.220 0.675 0.939 0.573 0.459 62.472
    LLaMA 3.1-70B 0.504 3.005 0.949 0.707 0.352 33.248
    LLaMA 3.1-8B 0.356 0.952 0.882 0.860 0.517 29.713
    Summscreen GPT-4o-mini 5.251 0.954 0.794 0.331 0.311 68.707
    LLaMA 3.1-70B 4.023 1.762 0.766 0.259 0.224 58.992
    LLaMA 3.1-8B 2.290 0.736 0.843 0.475 0.390 42.404
    Table 11: Performance on the 10% summary expansion task per dataset.
@@
    LLaMA 3.1-70B 1.449 0.782 0.814 0.500 0.386 41.466 3.800
    LLaMA 3.1-8B 1.495 0.828 0.851 0.568 0.418 36.493 4.480
    Table 12: Performance on the summary expansion task by a factor of 5, given the 20% summary input.
    The results of the summarization task of different levels are detailed in tables 1314 and 15. According to human evaluation, the best performing result in Table 13 is with GPT-4o-mini in the wikipedia domain and in Table 14 is with the same model in the legal domain (LexGlue). For table 15, the best results are with human output in the conversational domain (Summscreen).
@@
    LLaMA 3.1-70B 0.417 0.491 0.963 0.831 0.660 0.480 0.489 60.241 6.440
    LLaMA 3.1-8B 0.402 0.474 0.968 0.839 0.642 0.476 0.490 60.415 7.080
    Table 14: Performance on the 10% summarization task per dataset
@@
    LLaMA 3.1-70B 0.454 0.589 0.957 0.834 0.615 0.441 0.452 54.075 6.600
    LLaMA 3.1-8B 0.469 0.642 0.960 0.847 0.601 0.456 0.444 54.142 6.760
    Table 15: Performance on the 20% summarization task per dataset.